Watch out here comes a Fax!

to: You!

from: MyMame
address: MyStreet

MyTown MyCountry

phone: MyPhoneNum
fax: MyFaxNum

Date: March 5, 1996

Fax: 6 pages

Note: This message was sent using a fax-modem. When replying to this fax, phone first, so the fax-modem and PC can be switched on, and prepared for fax-reception. It is also possible to reach me at the following e-mail addresses.

jbrhebergen@et.tudedv.tudelft.nl or rheberg@morra.et.tudelft.nl

the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times! the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times! the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times! the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times! the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times! the quick brown dog jumped over the lazy fox 0 1 2 3.1459 4 5 6 7 8 9 times!